# Wavefront Curvature and Transverse Atomic Motion in Time-Resolved Atom Interferometry: Impact and Mitigation

### v1

@@ I Introduction
Light-pulse atom interferometry [1, 2] exploits interference between atomic wave packets travelling along spatially separated paths. Laser pulses coherently split, redirect, and recombine the wave packets, enabling precision measurements of external forces and fields. Atom interferometers have been successfully used in diverse applications, including inertial sensing [3, 4, 5], determinations of the fine-structure constant [6, 7, 8], measurements of Newton’s gravitational constant [9], and tests of the equivalence principle [10, 11]. See [12] for a comprehensive review.
More recently, single-photon atom interferometers based on narrow optical transitions in alkaline-earth-like atoms have been proposed for groundbreaking applications in fundamental physics [13], such as gravitational-wave detection [14, 15, 16] and ultra-light dark-matter searches [17, 18, 16]. These applications are examples of time-resolved measurements, where the interferometric phase is measured repeatedly to capture dynamic signals. Time-resolved atom interferometry introduces unique challenges, as it demands unprecedented stability and control of systematic effects to achieve the desired sensitivity.
One of the most significant sources of systematic error in atom interferometry arises from laser wavefront aberrations, which cause the interferometric phase to depend on transverse atomic motion. While the impact of wavefront aberrations has been extensively studied in the context of time-integrated atom interferometry [19, 20, 21, 22, 23], where, for example, they have been identified as the main systematic in measurements of the fine-structure constant [6, 7, 8], their effects on time-resolved measurements have not been studied in detail until now. This gap in understanding is critical, as next-generation, large-scale terrestrial atom interferometry experiments, such as AION [24], MAGIS-100 [25], MIGA [26], ELGAR [27] and ZAIGA [28], aim to achieve unprecedented sensitivity levels using large momentum transfer (LMT) techniques, which amplify the number of atom-light interactions, thereby increasing the influence of wavefront aberrations. This effect is also expected to impact future space-based atom interferometers [15].
@@ II Semi-Classical Modelling of the atom interferometer
We present the framework used to model the impact of Gaussian wavefront curvature on the phase shift measured by an atom interferometer. Our focus is on atom interferometers based on inelastic diffraction, where the internal state of the atoms is modified by the atom-light interactions [29], and where momentum is transferred via single-photon transitions. For simplicity, we also restrict our attention to two-level systems.
@@
We use a semi-classical (or ray atom optics) model [30, 31], in which atoms are treated as point-like particles following classical trajectories, while their internal states evolve quantum mechanically as a two-level system [31]. Each atom thus follows two classical trajectories (upper and lower arms) during the interferometric sequence, along which a complex phase is accumulated. We ignore the effect of parasitic paths [32]. For a given atom characterised by initial coordinates $(\bm{x}_{0},\bm{v}_{0})$, the total interferometric phase is
@@
where the quantities on the right-hand side are evaluated along classical trajectories defined by $(\bm{x}_{0},\bm{v}_{0})$. Here, $\varphi_{\text{tot},u}$ and $\varphi_{\text{tot},l}$ refer to the total phase accumulated along the upper and lower arm trajectories, respectively. The second term accounts for the phase difference due to spatial separation between trajectories at detection time [33].
This semi-classical approach has been successfully applied to quantify the effects of wavefront aberrations in atom interferometers [23, 34, 7, 35] and forms the basis of our analysis.
@@ II.1 Phase evolution along a single trajectory
where $\omega_{0}$ is the resonant frequency of the two-level system and $S_{\text{cl}}$ is the action evaluated along the classical trajectory between initial and final positions [36, 37].
@@
In the semi-classical model, we assume that the atom is sufficiently localized relative to $\phi(\bm{x})$ so that, at the $i$-th interaction, the laser field can be approximated locally by a plane wave with phase offset $\phi(\bm{x}_{i})$, where $\bm{x}_{i}$ is the position of the atom at that interaction. This case can be treated analytically [38, 33]. Under this approximation, each laser interaction at time $t_{i}$ contributes a phase
@@
Note that we neglect here the momentum imparted by the transverse phase gradient of the laser beam, which is typically much smaller than the momentum imparted by the longitudinal phase gradient [23].
@@ II.2 Phase difference in a LMT Mach-Zehnder interferometer
For a perfectly symmetric MZ sequence in a uniform gravitational field, the separation-induced phase vanishes and the propagation phases of the two arms cancel, so that the residual interferometer phase is entirely due to the laser field [37]. We therefore only consider the laser phase contributions in the rest of this work.
@@
In single-photon clock atom interferometry based on 87Sr, which will be used in future large-scale experiments like AION [24] and MAGIS-100 [25], pulse durations $\tau$ are $\sim$1 ms due to the narrow linewidth of the clock transition. For large $n$, the total duration of an LMT pulse block becomes significant, such that the space-time area enclosed by the interferometer reduces from $n\hbar kT^{2}/m$ to $n\hbar kT(T-n\tau)/m$ [38]. To account for this effect, we substitute
@@ II.3 Phase shift due to transverse motion in a Gaussian beam
Under the paraxial approximation, the phase of a Gaussian beam with wavelength $\lambda$ propagating along the +z+z-direction can be written as [39]
@@
In what follows, we neglect the Gouy-phase term $\psi(z)$ in Eq. (13). While the Gouy phase can introduce significant systematic shifts in time-integrated measurements (e.g., in determinations of the fine-structure constant [6, 7, 8]) its contribution to shot-to-shot variations in the interferometric phase is several orders of magnitude smaller than the phase shifts induced by wavefront curvature. More generally, we focus on transverse effects because, whether arising through the curvature term $R(z)$ or through $\psi(z)$, longitudinal contributions remain far weaker. This approximation is justified in Appendix A.
@@ II.4 Effective phase perturbation in a position-averaged measurement
where the equation in the first line uses the conservation of phase-space volume under Hamiltonian evolution [40], the second line uses the chain rule of probability, and the third line substitutes Eq. (21). We assume that all atoms are detected in either the ground or excited port, such that $N_{\text{tot}}=n_{g}+n_{e}$, and ignore effects due to parasitic paths or open interferometry ports [32].
@@
to extract estimates of the phase shift $\Delta\varphi$ and contrast $C$. We emphasize that here, $\Delta\varphi$ is a free parameter of the model (not to be confused with the phase shift in absence of wavefront aberrations $\Delta\varphi_{0}$, or $\Delta\varphi(\bm{x}_{0},\bm{v}_{0})$). Note that the model is not invertible as-is because $C$ and $\Delta\varphi$ are not linearly independent. In practice, $\Delta\varphi_{0}$ is scanned in a controlled manner to generate a set of measurement points, which are then jointly fit to Eq. (25) (see example in Ref. [37]).
@@
where $\bm{\mu}_{x_{0}}={(\mu_{x_{0}},0,\mu_{z_{0}})}^{T}$ and $\bm{\mu}_{v_{0}}={(\mu_{v_{x0}},0,\mu_{v_{z_{0}}})}^{T}$, and $\mathcal{N}(\bm{\mu},\bm{\Sigma})$ denotes a multivariate normal distribution with mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$ [41]. Although written in vector form for completeness, only the $x$-components are treated as varying parameters in what follows. Shot-to-shot variations in $\mu_{x_{0}}$ and $\mu_{v_{x0}}$ are thus encoded in $\bm{\theta}=(\mu_{x_{0}},\mu_{v_{x0}})$, while $\mu_{z_{0}}$ and $\mu_{v_{z0}}$ are held fixed. In practice, atom number and temperature fluctuations can also cause the spreads $\sigma_{x_{0}}$ and $\sigma_{v_{0}}$ to vary between shots. We discuss this case in section III.3.
@@ III Wavefront-curvature noise in long-baseline atom gradiometers
Atom gradiometers operate by simultaneously measuring the phase shift in two vertically-separated atom interferometers operated by a common laser beam. By taking differential measurements, they suppress common-mode noise sources such as laser phase noise [42], enabling time-resolved applications that are otherwise limited. As such, they are expected to serve as the workhorse for next-generation precision instruments, including gravitational wave detectors and dark matter searches [24, 25]. The sensitivity of atom gradiometers to science signals typically scales with the vertical separation, or baseline, between the two atom interferometers ($L$ in Fig. 2). This scaling has motivated the proposal of several 100-meter-scale experiments, such as MAGIS-100, which is already under construction, and AION-100. Larger, km-scale follow-up experiments have also been proposed [25, 24].
@@ III.2.2 Impact on LMT $\pi$-pulse efficiencies
This is because both interferometers are driven by the same beam, so a single pulse duration must provide efficient population transfers at both locations. This constrains the allowable variation of the Rabi frequency along the baseline. Since the rate of change of $w(z)$ is zero at the focus and becomes constant at $z\to\pm\infty$, placing the focus at $z=0$, or equivalently ensuring that both interferometers lie within the slowly varying region of the beam waist (small $L/z_{R}$), satisfies the efficiency condition. Alternatively, pulses that are robust to pulse-area errors, such as composite pulses [43], could be used to mitigate these limitations.
@@ III.3 Transverse motion requirements for next-generation long-baseline gradiometers
Achieving phase noise levels of $10^{-5}$ rad, the target for future detectors such as AION [24] and MAGIS-100 [25], imposes stringent requirements on transverse motion stability. For a 100-m instrument with $w_{0}=1$ cm and $n=1001$, operation in the HEHN configuration requires transverse COM fluctuations to be stabilized at the few-micron level in position and below the micron-per-second level in velocity. In the LELN configuration, the corresponding stability requirements relax to the tens-of-microns scale in position and to just under ten microns per second in velocity. These values reflect approximately an order-of-magnitude difference between the two limits, consistent with the expected trade-off between efficiency and curvature noise suppression observed in Fig. 5.
By comparison, state-of-the-art atom interferometers have achieved COM position and velocities at the tens-of-microns and tens-of-microns-per-second levels, respectively [44, 45]. Thus, meeting the HEHN requirements will require roughly a tenfold improvement over current experimental performance, whereas the LELN stability targets are already within reach, provided composite or robust pulse techniques [43] are employed to mitigate the reduced efficiency.
Note that, however, reaching these targets could be challenging for fermionic isotopes such as 87Sr, proposed for use in both AION and MAGIS-100, which will necessitate extremely weak trapping potentials. For instance, obtaining a trapped Fermi gas with $N=10^{6}$ atoms, while maintaining a characteristic position radius $x_{F}\sim\sqrt{\hbar N^{1/3}/m\omega}\sim$1 mm and velocity radius $v_{F}\sim\sqrt{\hbar N^{1/3}\omega/m}\lesssim$1 mms-1, requires significantly weaker trap frequencies on the order of $\omega\sim 2\pi\times(0.1\,\text{Hz}-1\,\text{Hz})$ [46].
@@
In addition, terrestrial atom interferometers are subject to phase shifts induced by the Coriolis force due to Earth’s rotation. This effect is linearly sensitive to initial transverse velocity fluctuations [47], and typically dominates over the curvature-induced phase noise discussed here, which depends quadratically on the initial position and velocity. Consequently, Coriolis-induced noise would generally impose even stricter requirements on transverse velocity stability than those derived above. However, future long-baseline atom interferometry applications, whether time-resolved or time-integrated, are expected to implement established Coriolis compensation techniques [48, 49]. The influence of residual couplings due to imperfect compensation could be the subject of future investigation.
@@ IV Mitigation strategy: position-resolved phase-shift readout
In the previous section, we assumed that all spatial information about the atom cloud was integrated over. Position-resolved phase-shift readout, however, has been identified as a promising mitigation strategy for systematic effects caused by wavefront aberrations in atom interferometry [25, 50]. The non-uniform wavefront makes the interferometric phase sensitive to the atoms’ transverse position and velocity, resulting in phase noise when these quantities are not measured. While position-resolved measurements directly provide position information, velocity can also be inferred indirectly through suitable analysis of the position-resolved phase. This principle is, for example, well established in point-source atom interferometry, where ballistic expansion creates a deterministic correlation between an atom’s final position and its initial velocity, allowing velocity-dependent phase shifts to be mapped onto the spatial distribution of the cloud at detection time [47, 51, 52].

### v2

@@ I Introduction
Light-pulse atom interferometry [YOUNG1997363, Borde1989] exploits interference between atomic wave packets travelling along spatially separated paths. Laser pulses coherently split, redirect, and recombine the wave packets, enabling precision measurements of external forces and fields. Atom interferometers have been successfully used in diverse applications, including inertial sensing [peters1999, PhysRevLett.78.2046, PhysRevA.65.033608], determinations of the fine-structure constant [Bouchendira2010, Parker, Morel2020], measurements of Newton’s gravitational constant [Rosi_2014], and tests of the equivalence principle [PhysRevA.88.043615, PhysRevLett.113.023005]. See [abend2020atominterferometryapplications] for a comprehensive review.
More recently, single-photon atom interferometers based on narrow optical transitions in alkaline-earth-like atoms have been proposed for groundbreaking applications in fundamental physics [Buchmueller03042023], such as gravitational-wave detection [PhysRevLett.110.171102, Hogan_2011, Badurina_2021] and ultra-light dark-matter searches [PhysRevLett.115.011802, PhysRevLett.117.261301, Badurina_2021]. These applications are examples of time-resolved measurements, where the interferometric phase is measured repeatedly to capture dynamic signals. Time-resolved atom interferometry introduces unique challenges, as it demands unprecedented stability and control of systematic effects to achieve the desired sensitivity.
One of the most significant sources of systematic error in atom interferometry arises from laser wavefront aberrations, which cause the interferometric phase to depend on transverse atomic motion. While the impact of wavefront aberrations has been extensively studied in the context of time-integrated atom interferometry [Wicht2005, Fils2005, Karcher2018, Louchet-Chauvet2011, Pagot2024], where, for example, they have been identified as the main systematic in measurements of the fine-structure constant [Bouchendira2010, Parker, Morel2020], their effects on time-resolved measurements have not been studied in detail until now. This gap in understanding is critical, as next-generation, large-scale terrestrial atom interferometry experiments, such as AION [Badurina_2020], MAGIS-100 [Abe_2021], MIGA [canuel2022gravityantennabasedquantum], ELGAR [canuel2020technologieselgarlargescale] and ZAIGA [Zhao_2022], aim to achieve unprecedented sensitivity levels using large momentum transfer (LMT) techniques, which amplify the number of atom-light interactions, thereby increasing the influence of wavefront aberrations. This effect is also expected to impact future space-based atom interferometers [Hogan_2011].
@@ II Semi-Classical Modelling of the atom interferometer
We present the framework used to model the impact of Gaussian wavefront curvature on the phase shift measured by an atom interferometer. Our focus is on atom interferometers based on inelastic diffraction, where the internal state of the atoms is modified by the atom-light interactions [Miffre2006], and where momentum is transferred via single-photon transitions. For simplicity, we also restrict our attention to two-level systems.
@@
We use a semi-classical (or ray atom optics) model [MeystreTextBook, Moore1997], in which atoms are treated as point-like particles following classical trajectories, while their internal states evolve quantum mechanically as a two-level system [Moore1997]. Each atom thus follows two classical trajectories (upper and lower arms) during the interferometric sequence, along which a complex phase is accumulated. We ignore the effect of parasitic paths [PhysRevLett.131.033602]. For a given atom characterised by initial coordinates $(\bm{x}_{0},\bm{v}_{0})$, the total interferometric phase is
@@
where the quantities on the right-hand side are evaluated along classical trajectories defined by $(\bm{x}_{0},\bm{v}_{0})$. Here, $\varphi_{\text{tot},u}$ and $\varphi_{\text{tot},l}$ refer to the total phase accumulated along the upper and lower arm trajectories, respectively. The second term accounts for the phase difference due to spatial separation between trajectories at detection time [hogan2008lightpulseatominterferometry].
This semi-classical approach has been successfully applied to quantify the effects of wavefront aberrations in atom interferometers [Pagot2024, Schkolnik_2015, Parker, Bade2018] and forms the basis of our analysis.
@@ II.1 Phase evolution along a single trajectory
where $\omega_{0}$ is the resonant frequency of the two-level system and $S_{\text{cl}}$ is the action evaluated along the classical trajectory between initial and final positions [Storey1994, PhysRevLett.67.181].
@@
In the semi-classical model, we assume that the atom is sufficiently localized relative to $\phi(\bm{x})$ so that, at the $i$-th interaction, the laser field can be approximated locally by a plane wave with phase offset $\phi(\bm{x}_{i})$, where $\bm{x}_{i}$ is the position of the atom at that interaction. This case can be treated analytically [antoine2006, hogan2008lightpulseatominterferometry]. Under this approximation, each laser interaction at time $t_{i}$ contributes a phase
@@
Note that we neglect here the momentum imparted by the transverse phase gradient of the laser beam, which is typically much smaller than the momentum imparted by the longitudinal phase gradient [Pagot2024].
@@ II.2 Phase difference in a LMT Mach-Zehnder interferometer
For a perfectly symmetric MZ sequence in a uniform gravitational field, the separation-induced phase vanishes and the propagation phases of the two arms cancel, so that the residual interferometer phase is entirely due to the laser field [PhysRevLett.67.181]. We therefore only consider the laser phase contributions in the rest of this work.
@@
In single-photon clock atom interferometry based on 87Sr, which will be used in future large-scale experiments like AION [Badurina_2020] and MAGIS-100 [Abe_2021], pulse durations $\tau$ are $\sim$1 ms due to the narrow linewidth of the clock transition. For large $n$, the total duration of an LMT pulse block becomes significant, such that the space-time area enclosed by the interferometer reduces from $n\hbar kT^{2}/m$ to $n\hbar kT(T-n\tau)/m$ [antoine2006]. To account for this effect, we substitute
@@ II.3 Phase shift due to transverse motion in a Gaussian beam
Under the paraxial approximation, the phase of a Gaussian beam with wavelength $\lambda$ propagating along the +z+z-direction can be written as [Svelto2010]
@@
In what follows, we neglect the Gouy-phase term $\psi(z)$ in Eq. (13). While the Gouy phase can introduce significant systematic shifts in time-integrated measurements (e.g., in determinations of the fine-structure constant [Bouchendira2010, Parker, Morel2020]) its contribution to shot-to-shot variations in the interferometric phase is several orders of magnitude smaller than the phase shifts induced by wavefront curvature. More generally, we focus on transverse effects because, whether arising through the curvature term $R(z)$ or through $\psi(z)$, longitudinal contributions remain far weaker. This approximation is justified in Appendix A.
@@ II.4 Effective phase perturbation in a position-averaged measurement
where the equation in the first line uses the conservation of phase-space volume under Hamiltonian evolution [Tolman1938], the second line uses the chain rule of probability, and the third line substitutes Eq. (21). We assume that all atoms are detected in either the ground or excited port, such that $N_{\text{tot}}=n_{g}+n_{e}$, and ignore effects due to parasitic paths or open interferometry ports [PhysRevLett.131.033602].
@@
to extract estimates of the phase shift $\Delta\varphi$ and contrast $C$. We emphasize that here, $\Delta\varphi$ is a free parameter of the model (not to be confused with the phase shift in absence of wavefront aberrations $\Delta\varphi_{0}$, or $\Delta\varphi(\bm{x}_{0},\bm{v}_{0})$). Note that the model is not invertible as-is because $C$ and $\Delta\varphi$ are not linearly independent. In practice, $\Delta\varphi_{0}$ is scanned in a controlled manner to generate a set of measurement points, which are then jointly fit to Eq. (25) (see example in Ref. [PhysRevLett.67.181]).
@@
where $\bm{\mu}_{x_{0}}={(\mu_{x_{0}},0,\mu_{z_{0}})}^{T}$ and $\bm{\mu}_{v_{0}}={(\mu_{v_{x0}},0,\mu_{v_{z_{0}}})}^{T}$, and $\mathcal{N}(\bm{\mu},\bm{\Sigma})$ denotes a multivariate normal distribution with mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$ [Gut1995]. Although written in vector form for completeness, only the $x$-components are treated as varying parameters in what follows. Shot-to-shot variations in $\mu_{x_{0}}$ and $\mu_{v_{x0}}$ are thus encoded in $\bm{\theta}=(\mu_{x_{0}},\mu_{v_{x0}})$, while $\mu_{z_{0}}$ and $\mu_{v_{z0}}$ are held fixed. In practice, atom number and temperature fluctuations can also cause the spreads $\sigma_{x_{0}}$ and $\sigma_{v_{0}}$ to vary between shots. We discuss this case in section III.3.
@@ III Wavefront-curvature noise in long-baseline atom gradiometers
Atom gradiometers operate by simultaneously measuring the phase shift in two vertically-separated atom interferometers operated by a common laser beam. By taking differential measurements, they suppress common-mode noise sources such as laser phase noise [baynham2025prototypeatominterferometerdetect], enabling time-resolved applications that are otherwise limited. As such, they are expected to serve as the workhorse for next-generation precision instruments, including gravitational wave detectors and dark matter searches [Badurina_2020, Abe_2021]. The sensitivity of atom gradiometers to science signals typically scales with the vertical separation, or baseline, between the two atom interferometers ($L$ in Fig. 2). This scaling has motivated the proposal of several 100-meter-scale experiments, such as MAGIS-100, which is already under construction, and AION-100. Larger, km-scale follow-up experiments have also been proposed [Abe_2021, Badurina_2020].
@@ III.2.2 Impact on LMT $\pi$-pulse efficiencies
This is because both interferometers are driven by the same beam, so a single pulse duration must provide efficient population transfers at both locations. This constrains the allowable variation of the Rabi frequency along the baseline. Since the rate of change of $w(z)$ is zero at the focus and becomes constant at $z\to\pm\infty$, placing the focus at $z=0$, or equivalently ensuring that both interferometers lie within the slowly varying region of the beam waist (small $L/z_{R}$), satisfies the efficiency condition. Alternatively, pulses that are robust to pulse-area errors, such as composite pulses [Genov2014], could be used to mitigate these limitations.
@@ III.3 Transverse motion requirements for next-generation long-baseline gradiometers
Achieving phase noise levels of $10^{-5}$ rad, the target for future detectors such as AION [Badurina_2020] and MAGIS-100 [Abe_2021], imposes stringent requirements on transverse motion stability. For a 100-m instrument with $w_{0}=1$ cm and $n=1001$, operation in the HEHN configuration requires transverse COM fluctuations to be stabilized at the few-micron level in position and below the micron-per-second level in velocity. In the LELN configuration, the corresponding stability requirements relax to the tens-of-microns scale in position and to just under ten microns per second in velocity. These values reflect approximately an order-of-magnitude difference between the two limits, consistent with the expected trade-off between efficiency and curvature noise suppression observed in Fig. 5.
By comparison, state-of-the-art atom interferometers have achieved COM position and velocities at the tens-of-microns and tens-of-microns-per-second levels, respectively [PhysRevA.101.033606, Asenbaum_2020]. Thus, meeting the HEHN requirements will require roughly a tenfold improvement over current experimental performance, whereas the LELN stability targets are already within reach, provided composite or robust pulse techniques [Genov2014] are employed to mitigate the reduced efficiency.
Note that, however, reaching these targets could be challenging for fermionic isotopes such as 87Sr, proposed for use in both AION and MAGIS-100, which will necessitate extremely weak trapping potentials. For instance, obtaining a trapped Fermi gas with $N=10^{6}$ atoms, while maintaining a characteristic position radius $x_{F}\sim\sqrt{\hbar N^{1/3}/m\omega}\sim$1 mm and velocity radius $v_{F}\sim\sqrt{\hbar N^{1/3}\omega/m}\lesssim$1 mms-1, requires significantly weaker trap frequencies on the order of $\omega\sim 2\pi\times(0.1\,\text{Hz}-1\,\text{Hz})$ [Dean2018].
@@
In addition, terrestrial atom interferometers are subject to phase shifts induced by the Coriolis force due to Earth’s rotation. This effect is linearly sensitive to initial transverse velocity fluctuations [PhysRevLett.111.083001], and typically dominates over the curvature-induced phase noise discussed here, which depends quadratically on the initial position and velocity. Consequently, Coriolis-induced noise would generally impose even stricter requirements on transverse velocity stability than those derived above. However, future long-baseline atom interferometry applications, whether time-resolved or time-integrated, are expected to implement established Coriolis compensation techniques [Lan2012, glick2023coriolisforcecompensationlaser]. The influence of residual couplings due to imperfect compensation could be the subject of future investigation.
@@ IV Mitigation strategy: position-resolved phase-shift readout
In the previous section, we assumed that all spatial information about the atom cloud was integrated over. Position-resolved phase-shift readout, however, has been identified as a promising mitigation strategy for systematic effects caused by wavefront aberrations in atom interferometry [Abe_2021, seckmeyer2024principalcomponentanalysisspatial]. The non-uniform wavefront makes the interferometric phase sensitive to the atoms’ transverse position and velocity, resulting in phase noise when these quantities are not measured. While position-resolved measurements directly provide position information, velocity can also be inferred indirectly through suitable analysis of the position-resolved phase. This principle is, for example, well established in point-source atom interferometry, where ballistic expansion creates a deterministic correlation between an atom’s final position and its initial velocity, allowing velocity-dependent phase shifts to be mapped onto the spatial distribution of the cloud at detection time [PhysRevLett.111.083001, Chen2020, PhysRevA.102.013326].